10. Show the ratio of the Compton wavelength to the de Broglie wavelength for a relativistic electron with Energy E is given by 소 11. Non-relativistic electrons and protons are accelerated from rest though the same potential difference Δ V, show the ratio of their de Broglie wavelengths is given by: 弖
Answer #1

10) The Compton wavelength is

\lambda _c=\frac{h}{m_ec}

The de Broglie wavelength is \lambda =\frac{h}{p} . By taking the ratio of the Compton wavelength to the de Broglie wavelength and square it

\left ( \frac{\lambda _c}{\lambda } \right )^2=\frac{p^2}{(m_ec)^2}

the momentum for a slowly-moving or rapidly moving object is described by

p^2=\frac{E^2-m_e^2c^4}{c^2}

Substituting and simplifying

\left ( \frac{\lambda_c}{\lambda } \right )^2=\frac{(E^2-m_e^2c^4)}{(m_ec^2)^2}=\left (\frac{E^2}{m_ec^2} \right )^2-1

\frac{\lambda_c}{\lambda }=\sqrt{\left (\frac{E^2}{E_o^2} \right )^2-1}